OpenStudy (anonymous):

Is r√r - 3r√r + 4r√r equal to 2r or 2√r? I know it has to be one of the two because 1 - 3 + 4 = 2, but I'm not sure if you are supposed to keep the radial sign or not.

hartnn (hartnn):

factor out r√r, so, r√r(1-3+4) = ... ?

OpenStudy (anonymous):

it's just r once you've done that, right?

OpenStudy (anonymous):

well, 2r I mean

hartnn (hartnn):

whats 1-3+4=...? then multiply that with r√r and its none of these : 2r or 2√r

OpenStudy (anonymous):

Then wouldn't that be 4r? And then the square of 4 is 2. The only other thing I can think of is that it could be 2r√r once it's been multiplied.

hartnn (hartnn):

' it could be 2r√r once it's been multiplied. ' <----it is. r√r(1-3+4) = r√r(2) = 2r√r.

OpenStudy (anonymous):

Ahh, sorry. That's right! I'm not sure what I was thinking. Thank you.

hartnn (hartnn):

welcome ^_^